# Analysis of Active/Inactive Patterns in the NHANES Data using Generalized Multilevel Functional Principal Component Analysis

### v2

@@ 3 Methods
3.1 Implementation
@@
The code below (Listing. 1) provides the implementation of Steps 1-3 of the algorithm for binary data, while the code for Step 4 is shown in the supplementary materials. The code is short and concise, and allows for easy implementation and modification.
@@
3.2 Practical considerations
Choosing the bin width. In practice, if the bin width is too small, there may not be enough variation between subjects or between visits among the same subject to fit the local GLMM model in certain bins; convergence or singularity issues can arise. The problem is exacerbated when the number of visits per subject is small (e.g., $2$-$5$). At the other extreme when bins are too large, the curvature of the latent process may not be well captured. This is because each local GLMM model essentially estimates a mean (with fixed and random components) in each bin and for each function. If the functions exhibit substantial curvature in a large bin, then the model would not provide a good fit to the data. From a practical standpoint, we have found that one does not need to find the perfect bin width to obtain reasonable results. Indeed, there are many choices that provide similar results and a good strategy is to try a few different bin widths.
@@ 3.2 Practical considerations
Choosing the number of eigenfunctions. The most common approach for choosing the number of eigenfunctions is based on the cumulative percentage of variance explained. While this strategy is effective, it sometimes leads to selecting too many eigenfunctions. Eigenfunctions that rank lower do not contribute much to the variance explained, slow down Step 4, and are more challenging to interpret because they typically show more oscillation compared to higher-ranked eigenfunctions. To inform the choice of the eigenfunctions, we propose to use the sum of squared second-order differences (SSSOD), which is the discrete analog of the integrated squared second-order derivative (ISSOD). A higher SSSOD corresponds to more oscillations in an eigenfunction, which reduces its interpretability.
@@
4 Simulation study
In this section we simulate binary and Poisson multilevel functional data and evaluate how well the proposed method can recover the eigenfunctions, scores, and latent functional trajectories. For both binary and Poisson data, the latent continuous response $\eta_{ij}(s_{k})$ is generated from the following model:
$\eta_{ij}(s_{k})$ $=$ $b_{0}(s_{k})+\sum_{l=1}^{4}\xi_{il}\phi_{l}^{(1)}(s_{k})+\sum_{m=1}^{4}\zeta_{ijm}\phi_{m}^{(2)}(s_{k}),$
where $b_{0}(s_{k})$ denotes the intercept at time $s_{k}$, $\xi_{il}\sim N(0,\lambda_{l}^{(1)})$, $\zeta_{ijm}\sim N(0,\lambda_{m}^{(2})$, $\{s_{k}=\frac{k}{K}:k=0,1,...,K\}$, and $K$ is the number of sampling points. We set the true eigenvalues $\lambda_{l}^{(1)}=0.5^{l-1},l=1,2,3,4$ and $\lambda_{m}^{(2)}=0.5^{m-1},m=1,2,3,4$. For simulating eigenfunctions, we follow (Di et al.,, 2009) and consider two cases:
Case 1. Mutually orthogonal bases.
  1. Level 1:
  2. Level 2:
Case 2. Mutually nonorthogonal bases between the two levels.
  1. Level 1:
    same as Case 1.
  2. Level 2:
Multilevel binary functional data are generated as $Z_{ij}(s_{k})\sim\text{Bernoulli}\{\mu_{ij}(s_{k})\}$, where ${\rm logit}\{\mu_{ij}(s_{k})\}=\eta_{ij}(s_{k})$. Multilevel Poisson data are generated as $Z_{ij}(s_{k})\sim\text{Poisson}\{\mu_{ij}(s_{k})\}$, where $\log\{\mu_{ij}(s_{k})\}=\eta_{ij}(s_{k})$.
For Bernoulli we considered $32$ simulation scenarios and for Poisson we considered $26$ simulation scenarios. For each scenario we ran $20$ simulations. More precisely, in addition to varying the eigenfunctions ($2$ scenarios), we also examined the effect of sample size $I=50,100,200,500,1000$, number of visits $J=2,5,10$, and number of sampling points $K=100,200,500$. In the binary case the proportion of ones was controlled by the choice of the intercept function. We considered a constant intercept function $b_{0}(s)=b_{0}$ and varied $b_{0}=0,-1.5,-2.5,-3.5$, which corresponds to roughly $50$%, $30$%, $17$%, and $9$% of ones. When fitting models in Step 2 we have considered bins of a length corresponding to $w=2\%,5\%,10\%,15\%,20\%$ of the number of sampling points. In step 4, we set the number of post warm-up iterations to $1,000$. Running simulations for all combinations of parameter settings would have resulted in $1,800$ scenarios for Bernoulli and $450$ for Poisson regressions, an impossible task given our computational resources. Instead, we varied one parameter at a time while keeping other parameters fixed.
To evaluate the proposed method, we calculate the mean integrated squared error (MISE) of linear predictors across all subject-visit pairs
@@ 2. Level 2:
and the integrated squared error (ISE) of each eigenfunction:
[…]
for $l=1,2$ (two levels) and $m=1,2,3,4$.
4.1 Simulation results
Table 1 summarises the MISE of linear predictors for binary and Poisson data using case 1 and case 2 basis functions and different number of sampling points. For binary data, results show that as the number of sampling points increases, the proposed method achieves higher accuracy in recovering the latent linear predictors. For Poisson data, MISE decreases as $K$ increases from $100$ to $200$. At $K=500$, however, we found that the MCMC algorithm in Step 4 converged successfully for only half of the 20 simulated data sets. For this reason, we do not report the MISE at $K=500$ for Poisson data. Among the data sets where MCMC converged, the mean MISE at $K=500$ is less than half of the MISE at $K=200$. The non-convergence of MCMC may be caused by poor initialization or the sample size being too small relative to the number of sampling points. A more careful look into this issue is left to future work.
@@ 4.1 Simulation results
Table 1: MISE of linear predictors averaged over 20 simulations for different numbers of sampling points.
Simulation results for when the number of sampling points is fixed and the number of study participants, the number of visits, the bin width, or the fixed effect intercept increases are displayed in Table S.2 to S.8 in the supplementary materials. Indeed, as the sample size (number of subjects) increases, Table S.1 and S.2 show that both the ISE of eigenfunctions and the MISE of linear predictors generally go down. Table S.3 and S.4 show that a similar decrease is observed when the number of visits increases. As the bin width increases, however, Table S.5 shows that the ISE of eigenfunctions first goes down but then goes up. As explained in detail in Section 3.2, this pattern is expected: the decrease is driven by having more data included in each bin, and the increase is driven by using local models, which assumes a constant mean (with fixed and random components) in each bin, to approximate functions that can exhibit substantial curvature in a large bin. Finally, for binary data, Table S.7 and S.8 show that both ISE and MISE go down when the percentages of zeros and ones become more balanced (closer to $50$%). Results are similar for case 1 and case 2 basis functions.
4.2 Why Step 4 is necessary
4.3 Computation time
The computational bottleneck of GM-FPCA is in step 4, where subject-visit scores are estimated using the full data set. With $1,000$ post warm-up iterations, the median runtime ranges from half a day to 12 days depending on the scenario. The runtime of Step 4 for Poisson data at $K=500$ is excluded because of the non-convergence issue described in the previous section. See Table S.9 and S.10 from the supplementary materials for runtime results at different sample sizes and number of visits.
5 Application to NHANES accelerometry data
5.1 Data description
Since the goal of our analysis is to study mortality and its association with physical activity patterns, we excluded $9,149$ subjects under the age of 50, where age was ascertained at the time of their interview with NHANES. We further excluded 9 subjects whose mortality status is missing. The analytic sample for the GM-FPCA analysis thus contains $4,445$ subjects. With up to 7 days of accelerometry data per subject, we have $28,023$ days of wear and $1,440$ (the number of minutes in a day) observations per day. Table 3 summarizes the demographics and traditional mortality risk factors of the analytic sample. As mentioned before, we thresholded MIMS at $10.558$ (active if MIMS $>10.558$ and inactive otherwise) (Karas et al.,, 2022) to obtain binary active/inactive profiles.
5.2 Analysis results
The title for each eigenfunction subplot indicates the level, eigenfunction number, cumulative fraction of variance explained within the level, and sum of square second-order differences (SSSOD) for each function. For example, the panel in the first row and second column of Figure 5 is labeled “l1e2 (0.61 / 27.57)”. This indicates that this is the second eigenfunction (e2) at level one (l1), that the first two level one eigenfunctions (“l1e1” and “l1e2”) explain $61$% of the variability in level one, and that the SSSOD = $27.57$. We do not propose a rigorous analysis of the SSSOD measure at this time, but point out that it increases substantially with the observed complexity of the functions. For example, the SSSOD of eigenfunctions one and two are one or two orders of magnitude smaller compared to that of eigenfunctions five and six, and these differences track with the visual inspection of the plots. Indeed, the first three eigenfunctions are relatively easy to interpret because they have smaller variation and together they explain $77$% of the variation in level one. In contrast, the other eigenfunctions become increasingly difficult to interpret and together explain less than $23$% of the variability at this level.
After fitting the global model on each of the subsets, we obtain both level one and level two scores. We now investigate whether these scores are associated with mortality. More precisely, we investigate whether the subject-specific scores (level one) and the mean of subject-visit-specific scores (level two) are associated with time to death after accounting for known predictors of mortality. After further excluding 127 subjects who have missing values in at least one of the covariates (see Table 3 for the exact number of missingness for each covariate), we fit Cox proportional hazards models for the time to death outcome using three different sets of covariates. Model 1 only includes traditional risk factors for mortality such as age, gender, and BMI; Model 2 adds level one scores; Model 3 further adds the mean of level two scores. Table 4 displays the results from the three models. In addition to hazard ratios and 95% confidence intervals, we also show the concordance and the integrated Brier scores for all three models.
From model 2 and model 3, we see that the scores of the first three level one eigenfunctions are strongly predictive of mortality even after accounting for traditional risk factors. For example, from model 3 the hazard ratio for “l1e1” is 0.71 (95% CI [0.64, 0.77]), suggesting that subjects with higher scores on the level one eigenfunction 1 have lower risk of death when keeping other covariates in the model fixed. We have shown in the left panel of Figure 6 that subjects with higher “l1e1” scores tend to be more active during the day and less active during the night compared to those with lower “l1e1” scores, so the result agrees with our knowledge of what a healthy lifestyle should look like.
From model 3, we see that the first two mean level two scores are strongly predictive of mortality after accounting for both traditional risk factors and level one scores. To intuitively understand what level two eigenfunction 1 represents, we took two subjects whose mean level two eigenfunction 1 scores are at the extremes: one at the $1^{\text{st}}$ percentile and the other at the $99^{\text{th}}$ percentile. Then we smooth their binary activity data using the same approach as before (generalized additive models with cyclic cubic splines) and plot their smoothed activity profile from all seven days in the left panel of Figure 8 (each color represents a subject; each line represents a day). The figure shows that the subject whose mean level two eigenfunction 1 score is high (red, “l2e1-high”) has a less consistent daytime routine than the subject with a low score. (In fact, the subject shown in red has a less consistent nighttime routine as well, but because “l2e1” is close to 0 during the night, here we only focus on the daytime routine.) Since the hazard ratio for “l2e1” is 1.35 (95% CI [1.19, 1.54]), the result suggests that maintaining a consistent daytime routine is protective against death. While the conclusion agrees with common sense, we are unaware of any study that examined that quantified this association. We emphasize that this discovery would not have been possible with a single-level GFPCA analysis, which averages data over days within the same individual. With that said, we want to be transparent and acknowledge that the two subjects being plotted were not selected at random. Since the level two eigenfunction 1 only explains $10$% of the variability in level two, it does not have as strong a visual effect as top level one eigenfunctions do, so we chose two study participants whose differences are visually apparent. A similar plot is made for level two eigenfunction 2 and displayed in the right panel of Figure 8. We also tried adding the standard deviation of level two scores in the model, but they were not associated with mortality; results not reported.
Figure 8: Smoothed daily binary activity profile for two study participants whose level two eigenfunction 1 (left) and eigenfunction 2 (right) scores are at the $1^{\text{st}}$ percentile (blue) and the $99^{\text{th}}$ percentile (red). The estimated hazard ratios for “l2e1” and “l2e2” are 1.35 and 1.27, respectively, so participants in red have higher mortality risk.
@@ 5.2 Analysis results
Table 3: Summary of demographic variables and traditional risk factors for subjects over the age of 50 from the NHANES 2011-2014 accelerometry data set.

### v3

@@ 3 Methods
The R code for implementing the algorithm and a discussion of practical issues such as choosing the bin width and the number of eigenfunctions can be found in supplementary materials.
4 Simulation study
In this section we simulate binary and Poisson multilevel functional data and evaluate how well the proposed method can recover the eigenfunctions, scores, and latent functional trajectories. For both binary and Poisson data, the latent continuous response $\eta_{ijk}$ is generated from the following model:
$\eta_{ijk}$ $=$ $\beta_{0}(s_{k})+\sum_{l=1}^{4}\xi_{il}\phi_{l}^{(1)}(s_{k})+\sum_{m=1}^{4}\zeta_{ijm}\phi_{m}^{(2)}(s_{k}),$
where $\beta_{0}(s_{k})$ denotes the intercept at time $s_{k}$, $\xi_{il}\sim N(0,\lambda_{l}^{(1)})$, $\zeta_{ijm}\sim N(0,\lambda_{m}^{(2})$, $\{s_{k}=\frac{k}{K}:k=0,1,...,K\}$, and $K$ is the number of sampling points. We set the true eigenvalues $\lambda_{l}^{(1)}=0.5^{l-1},l=1,2,3,4$ and $\lambda_{m}^{(2)}=0.5^{m-1},m=1,2,3,4$. For simulating eigenfunctions, we follow (Di et al.,, 2009) and consider two cases:
Case 1. Mutually orthogonal bases.
  1. Level 1:
  2. Level 2:
Case 2. Mutually nonorthogonal bases between the two levels.
  1. Level 1:
    same as Case 1.
  2. Level 2:
Multilevel binary functional data are generated as $Z_{ijk}\sim\text{Bernoulli}\{\mu_{ijk}\}$, where ${\rm logit}\{\mu_{ijk}\}=\eta_{ijk}$. Multilevel Poisson data are generated as $Z_{ijk}\sim\text{Poisson}\{\mu_{ijk}\}$, where $\log\{\mu_{ijk}\}=\eta_{ijk}$.
For Bernoulli we considered $32$ simulation scenarios and for Poisson we considered $26$ simulation scenarios. For each scenario we ran $20$ simulations. More precisely, in addition to varying the eigenfunctions ($2$ scenarios), we also examined the effect of sample size $I=50,100,200,500,1000$, number of visits $J=2,5,10$, and number of sampling points $K=100,200,500$. In the binary case the proportion of ones was controlled by the choice of the intercept function. We considered a constant intercept function $\beta_{0}(s)=b_{0}$ and varied $b_{0}=0,-1.5,-2.5,-3.5$, which corresponds to roughly $50$%, $30$%, $17$%, and $9$% of ones. When fitting models in Step 2 we have considered bins of a length corresponding to $w=2\%,5\%,10\%,15\%,20\%$ of the number of sampling points. In step 4, we set the number of post warm-up iterations to $1,000$. Running simulations for all combinations of parameter settings would have resulted in $1,800$ scenarios for Bernoulli and $450$ for Poisson regressions, an impossible task given our computational resources. Instead, we varied one parameter at a time while keeping other parameters fixed.
To evaluate the proposed method, we calculate the mean squared error (MSE) of linear predictors across all subject-visit pairs
$\frac{1}{I\times J\times K}\sum_{i=1}^{I}\sum_{j=1}^{J}\sum_{k=1}^{K}[\widehat{\eta}_{ijk}-\eta_{ijk}]^{2},$
and the integrated squared error (ISE) of each eigenfunction:
[…]
for $l=1,2$ (two levels) and $m=1,2,3,4$.
4.1 Simulation results
Refer to captionRefer to caption
Refer to captionRefer to caption
Table 1 summarises the MSE of linear predictors for binary and Poisson data using case 1 and case 2 basis functions and different number of sampling points. For binary data, results show that as the number of sampling points increases, the proposed method achieves higher accuracy in recovering the latent linear predictors. For Poisson data, MSE decreases as $K$ increases from $100$ to $200$. At $K=500$, however, we found that the MCMC algorithm in Step 4 converged successfully for only half of the 20 simulated data sets. For this reason, we do not report the MSE at $K=500$ for Poisson data. Among the data sets where MCMC converged, the mean MSE at $K=500$ is less than half of the MSE at $K=200$. The non-convergence of MCMC may be caused by poor initialization or the sample size being too small relative to the number of sampling points. A more careful look into this issue is left to future work.
Table 1: MSE of linear predictors averaged over 20 simulations for different numbers of sampling points.
Simulation results for when the number of sampling points is fixed and the number of study participants, the number of visits, the bin width, or the fixed effect intercept increases are displayed in Table S.2 to S.8 in the supplementary materials. Indeed, as the sample size (number of subjects) increases, Table S.1 and S.2 show that both the ISE of eigenfunctions and the MSE of linear predictors generally go down. Table S.3 and S.4 show that a similar decrease is observed when the number of visits increases. As the bin width increases, however, Table S.5 shows that the ISE of eigenfunctions first goes down but then goes up. As explained in detail in supplementary materials Section S.2, this pattern is expected: the decrease is driven by having more data included in each bin, and the increase is driven by using local models, which assumes a constant mean (with fixed and random components) in each bin, to approximate functions that can exhibit substantial curvature in a large bin. Finally, for binary data, Table S.7 and S.8 show that both ISE and MSE go down when the percentages of zeros and ones become more balanced (closer to $50$%). Results are similar for case 1 and case 2 basis functions.
4.2 Why Step 4 is necessary
Refer to captionRefer to captionRefer to caption
4.3 Computation time
The computational bottleneck of GM-FPCA is in step 4, where subject-visit scores are estimated using the full data set. With $1,000$ post warm-up iterations, the median runtime ranges from half a day to 12 days depending on the scenario. The runtime of Step 4 for Poisson data at $K=500$ is excluded because of the non-convergence issue described in the previous section. See Table S.10 and S.11 from the supplementary materials for runtime results at different sample sizes and number of visits.
5 Application to NHANES accelerometry data
5.1 Data description
Since the goal of our analysis is to study mortality and its association with physical activity patterns, we excluded $9,149$ subjects under the age of 50, where age was ascertained at the time of their interview with NHANES. We further excluded 9 subjects whose mortality status is missing. The analytic sample for the GM-FPCA analysis thus contains $4,445$ subjects. With up to 7 days of accelerometry data per subject, we have $28,023$ days of wear and $1,440$ (the number of minutes in a day) observations per day. Table LABEL:supp-table:table1 from supplementary materials summarizes the demographics and traditional mortality risk factors of the analytic sample. As mentioned before, we thresholded MIMS at $10.558$ (active if MIMS $>10.558$ and inactive otherwise) (Karas et al.,, 2022) to obtain binary active/inactive profiles.
5.2 Analysis results
The title for each eigenfunction subplot indicates the level, eigenfunction number, cumulative fraction of variance explained within the level, and sum of square second-order differences (SSSOD, see supplementary materials Section S.2 for the definition) for each function. For example, the panel in the first row and second column of Figure 5 is labeled “l1e2 (0.61 / 27.57)”. This indicates that this is the second eigenfunction (e2) at level one (l1), that the first two level one eigenfunctions (“l1e1” and “l1e2”) explain $61$% of the variability in level one, and that the SSSOD = $27.57$. We do not propose a rigorous analysis of the SSSOD measure at this time, but point out that it increases substantially with the observed complexity of the functions. For example, the SSSOD of eigenfunctions one and two are one or two orders of magnitude smaller compared to that of eigenfunctions five and six, and these differences track with the visual inspection of the plots. Indeed, the first three eigenfunctions are relatively easy to interpret because they have smaller variation and together they explain $77$% of the variation in level one. In contrast, the other eigenfunctions become increasingly difficult to interpret and together explain less than $23$% of the variability at this level.
Refer to captionRefer to captionRefer to caption
After fitting the global model on each of the subsets, we obtain score estimates. We now investigate whether the subject-specific scores (level one) are associated with time to death after accounting for known predictors of mortality. After further excluding 127 subjects who have missing values in at least one of the covariates (see supplementary materials Table LABEL:supp-table:table1 for the exact number of missingness for each covariate), we fit Cox proportional hazards models for the time to death outcome using two sets of covariates. Model 1 only includes traditional risk factors for mortality such as age, gender, and BMI; Model 2 further adds level one scores. Table 3 displays the results from the two models. In addition to hazard ratios and 95% confidence intervals, we also show the concordance and the integrated Brier scores for both models.
From model 2, we see that the scores of the first three level one eigenfunctions are strongly predictive of mortality even after accounting for traditional risk factors. For example, the hazard ratio for “l1e1” is 0.85 (95% CI [0.79, 0.92]), suggesting that subjects with higher scores on the level one eigenfunction 1 have lower risk of death when keeping other covariates in the model fixed. We have shown in the left panel of Figure 6 that subjects with higher “l1e1” scores tend to be more active during the day and less active during the night compared to those with lower “l1e1” scores, so the result agrees with our knowledge of what a healthy lifestyle should look like.
Table 3: Analysis of time to mortality in NHANES using Cox proportional hazards models with two sets of covariates.
Hazard Ratio (95% CI)
Model 1 Model 2
Age 1.09∗∗∗ (1.08, 1.10) 1.09∗∗∗ (1.08, 1.10)
Gender: Female 0.72∗∗∗ (0.61, 0.84) 0.78∗∗ (0.66, 0.92)
BMI (Ref: Normal)
Underweight 1.33 (0.88, 2.03) 1.36 (0.90, 2.07)
Overweight 0.63∗∗∗ (0.51, 0.76) 0.62∗∗∗ (0.51, 0.75)
Obese 0.63∗∗∗ (0.51, 0.77) 0.59∗∗∗ (0.48, 0.72)
Race (Ref: Non-Hispanic White)
Non-Hispanic Black 0.80 (0.66, 0.97) 0.83 (0.68, 1.01)
Mexican American 0.63∗∗ (0.45, 0.88) 0.65 (0.47, 0.91)
Non-Hispanic Asian 0.50∗∗∗ (0.34, 0.72) 0.51∗∗∗ (0.36, 0.74)
Other Hispanic 0.62∗∗ (0.45, 0.86) 0.65 (0.47, 0.90)
Other 1.12 (0.66, 1.88) 1.13 (0.67, 1.91)
Overrall Health (Ref: Excellent)
Very good 0.89 (0.64, 1.24) 0.92 (0.66, 1.28)
Good 0.75 (0.60, 0.95) 0.77 (0.61, 0.97)
Fair 1.26 (1.05, 1.52) 1.28 (1.06, 1.54)
Poor 1.85∗∗∗ (1.41, 2.43) 1.92∗∗∗ (1.46, 2.53)
Diabetes (Ref: No)
Borderline 0.82 (0.53, 1.26) 0.83 (0.54, 1.28)
Yes 1.18 (0.99, 1.41) 1.18 (0.99, 1.40)
Arthritis: Yes 0.89 (0.75, 1.04) 0.88 (0.75, 1.04)
CHF: Yes 1.48∗∗∗ (1.18, 1.86) 1.43∗∗ (1.14, 1.80)
CHD: Yes 1.15 (0.91, 1.46) 1.17 (0.92, 1.48)
Heart Attack: Yes 1.20 (0.93, 1.53) 1.21 (0.95, 1.56)
Stroke: Yes 1.16 (0.93, 1.44) 1.11 (0.89, 1.38)
Cancer: Yes 1.28∗∗ (1.08, 1.52) 1.27∗∗ (1.07, 1.51)
Mobility Problem: Any difficulty 2.11∗∗∗ (1.76, 2.51) 2.06∗∗∗ (1.72, 2.45)
Alcohol Consumption (Ref: Never drinker)
Former drinker 1.26 (0.99, 1.61) 1.26 (0.98, 1.60)
Moderate drinker 0.83 (0.65, 1.07) 0.88 (0.69, 1.12)
Heavy drinker 1.25 (0.84, 1.85) 1.33 (0.90, 1.98)
Not available 1.10 (0.77, 1.57) 1.12 (0.78, 1.60)
Cigarette Smoking (Ref: Never)
Former 1.06 (0.89, 1.27) 1.08 (0.90, 1.28)
Current 1.30 (1.02, 1.65) 1.29 (1.01, 1.64)
l1e1 0.85∗∗∗ (0.79, 0.92)
l1e2 1.14 (1.03, 1.27)
l1e3 0.89∗∗ (0.81, 0.97)
l1e4 0.94 (0.81, 1.09)
l1e5 0.93 (0.78, 1.10)
Concordance 0.806 0.810
Integrated Brier score 0.068 0.066
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
6 Discussion
To summarize, we developed a novel statistical method called GM-FPCA for analyzing generalized (non-Gaussian) multilevel functional data, and confirmed activity patterns that are associated with higher mortality through an analysis of the binary NHANES 2011-2014 accelerometry data. These findings are easy to interpret and has the potential for wide public health impact. We believe that GM-FPCA would prove indispensable for researchers working with large wearable device data. Therefore, we make the code of GM-FPCA available both in the paper and as part of the R package fastGFPCA, available at https://github.com/julia-wrobel/fastGFPCA.
Supplementary Material
S.1 Implementation
@@ 6 Discussion
The code below (Listing. 1) provides the implementation of Steps 1-3 of the algorithm for binary data. The code is short and concise, and allows for easy implementation and modification.
@@
This code below (Listing. 2) provides the Stan code for implementing Step 4 of the algorithm. We only provide code for binary outcomes, but it can easily be modified for Poisson outcomes.
Listing 2: Stan code for Step 4 of the algorithm.
1functions {
3 return bernoulli_logit_lupmf(y | eta[start:end]);
4 }
5}
6data {
7 int<lower=0> N; // number of individuals
8 int<lower=0> J; // number of repeated functions within individuals
9 int<lower=0> S; // number of observations per function
10 int<lower=0> K; // number of level 1 eigenfunctions
11 int<lower=0> L; // number of level 2 eigenfunctions
12 int<lower=0> Nobs; // number of total observations
14 matrix[K,S+1] efuncs_l1; // level 1 eigen functions
15 matrix[L,S+1] efuncs_l2; // level 2 eigen functions
16 int<lower=1> grainsize;
17}
18parameters {
19 real beta_0; // global avereage
20 matrix[N, K] xil1; // level 1 random effects
21 matrix[N*J, L] xil2; // level 2 random effects
22
23 vector<lower=0>[K] sigmal1sq; // var of the l1 random intercept
24 vector<lower=0>[L] sigmal2sq; // var of the l2 random intercept
25}
26transformed parameters {
27 matrix[N, K] xil1_sc; // level 1 random effects
28 matrix[N*J, L] xil2_sc; // level 2 random effects
29 for(k in 1:K){
30 xil1_sc[,k] = xil1[,k] * sqrt(sigmal1sq[k]);
31 }
32 for(k in 1:L){
33 xil2_sc[,k] = xil2[,k] * sqrt(sigmal2sq[k]);
34 }
35}
36model {
37 int inx;
38 int inx_ij;
39 vector[Nobs] eta_vec;
40 matrix[N, S+1] eta_mat_i;
41 matrix[N*J, S+1] eta_mat_ij;
42
43 // priors
44 beta_0 ~ normal(0, 10);
45 to_vector(xil1) ~ normal(0,1);
46 to_vector(xil2) ~ normal(0,1);
47
48 sigmal1sq ~ inv_gamma(1, 1);
49 sigmal2sq ~ inv_gamma(1, 1);
50
51 eta_mat_i = xil1_sc * efuncs_l1;
52 eta_mat_ij = xil2_sc * efuncs_l2;
53 inx = 1;
54 for(i in 1:N){
55 for(j in 1:J){
56 eta_mat_ij[inx,] = eta_mat_ij[inx,] + eta_mat_i[i,];
57 inx = inx + 1;
58 }
59 }
60 eta_vec = to_vector((eta_mat_ij)’);
61
63}’
S.2 Practical considerations
S.2.1 Choosing the bin width
In practice, if the bin width is too small, there may not be enough variation between subjects or between visits among the same subject to fit the local GLMM model in certain bins; convergence or singularity issues can arise. The problem is exacerbated when the number of visits per subject is small (e.g., $2$-$5$). At the other extreme when bins are too large, the curvature of the latent process may not be well captured. This is because each local GLMM model essentially estimates a mean (with fixed and random components) in each bin and for each function. If the functions exhibit substantial curvature in a large bin, then the model would not provide a good fit to the data. From a practical standpoint, we have found that one does not need to find the perfect bin width to obtain reasonable results. Indeed, there are many choices that provide similar results and a good strategy is to try a few different bin widths.
@@
S.2.2 Choosing the number of eigenfunctions
The most common approach for choosing the number of eigenfunctions is based on the cumulative percentage of variance explained. While this strategy is effective, it sometimes leads to selecting too many eigenfunctions. Eigenfunctions that rank lower do not contribute much to the variance explained, slow down Step 4, and are more challenging to interpret because they typically show more oscillation compared to higher-ranked eigenfunctions. To inform the choice of the eigenfunctions, we propose to use the sum of squared second-order differences (SSSOD), which is the discrete analog of the integrated squared second-order derivative (ISSOD). A higher SSSOD corresponds to more oscillations in an eigenfunction, which reduces its interpretability.
@@
S.3 More on Step 4
S.3.1 Prior distribution
@@
S.3.2 Downsampling in Step 4
This section explores downsampling along the functional domain in Step 4 to speed up the estimation of scores. We emphasize that downsampling is not needed in Step 2 because fitting GLMMs in local bins is fully parallelizable; it is also not needed in Step 3 because MFPCA (Cui et al., 2022b, ) is highly scalable.
Refer to captionRefer to caption
S.4 More simulation results
S.4.1 Effect of Sample Size
In this experiment, we vary the sample size (number of study participants) while fixing other parameters at the following values: number of visits $J=10$, number of sampling points $K=100$, and percentage of data used for constructing local bins $w=5\%$. For binary data, we used fixed effect intercept $b_{0}=0$. Table S.1 and S.2 display the ISE of eigenfunctions and MSE of linear predictors for different data distributions and basis functions as a function of sample size $I=50,100,200,500,1000$. Results indicate that ISE decreases quickly with increased sample size and that linear predictors are estimated increasingly well when the number of study participants increases.
@@
Table S.2: MSE of linear predictors averaged over 20 repeats for different sample sizes.
S.4.2 Effect of the number of visits
In this experiment, we vary the number of visits while fixing other parameters at the following values: sample size $I=1000$, number of sampling points $K=100$, and percentage of data used for constructing local bins $w=5\%$. For binary data, we used fixed effect intercept $b_{0}=0$. Table S.3 and S.4 display the ISE of eigenfunctions and MSE of linear predictors for different data distributions and basis functions as a function of the number of visits $J=2,5,10$. Results indicate that both ISE and MSE are estimated increasingly well when the number of visits increases.
Table S.4: MSE of linear predictors averaged over 20 repeats for different numbers of visits.
S.4.3 Effect of bin width
In this experiment, we vary the bin width while fixing other parameters at the following values: sample size $I=1000$, number of visits $J=10$, and number of sampling points $K=100$. For binary data, we used fixed effect intercept $b_{0}=0$. Table S.5 and S.6 display the ISE of eigenfunctions and MSE of linear predictors for different data distributions and basis functions as a function of the bin width, expressed as a percentage of the number of total sampling points. We can see that with the exception of Poisson data using case 2 eigenfunctions, the MSE linear predictors first goes down and then goes up. Similar trend is observed for the ISE of many eigenfunctions too. To understand what happened to Poisson data using case 2 eigenfunctions, we further conducted experiments at $w=3\%$ and $4\%$. The MSE of linear predictors are $0.092$ and $0.093$, respectively, which suggest that $w=2\%$ is likely the best bin width for this scenario.
Table S.6: MSE of linear predictors averaged over 20 repeats for different bin widths.
S.4.4 Effect of the fixed-effect intercept
In this experiment, we focus on binary data and vary the fixed-effect intercept while fixing other parameters at the following values: sample size $I=1000$, number of visits $J=10$, number of sampling points $K=100$, and percentage of data used for constructing local bins $w=5\%$. Table S.7 and S.8 display the ISE of eigenfunctions and MSE of linear predictors for different data distributions and basis functions as a function of the intercept. Results indicate that both ISE and MSE are estimated increasingly well when the percentage of zeros and ones become more balanced (closer to $50$%).
Table S.8: MSE of linear predictors averaged over 20 repeats for different intercepts.
S.5 More NHANES data results
S.5.1 Summary of demographic variables and traditional risk factors
@@
Table S.9: Summary of demographic variables and traditional risk factors for subjects over the age of 50 from the NHANES 2011-2014 accelerometry data set.